OpenStudy (ashlyhall8):

Each blank should be filled with a single whole number greater than 1. Factor: 900 - 1 =

OpenStudy (mjdennis):

What have you tried so far?

OpenStudy (ashlyhall8):

Nothing i dont know how to do it

OpenStudy (welshfella):

this is the difference of 2 squares 900 = 30*30 and 1 = 1*1

OpenStudy (mjdennis):

OK, what subject are you studying? It looks like an algebra trick will solve it, and I want to see if you can find it.

OpenStudy (welshfella):

have you done the difference of 2 squares in algebra class?

OpenStudy (mjdennis):

@welshfella , you got this.

OpenStudy (welshfella):

@mjdennis you can have it I have to go right now

OpenStudy (mjdennis):

@ashlyhall8 , I have to go in 15 minutes, too. do we have you on track? Have you learned how to factor numbers that have the form \[(a ^{2} - b ^{2})\]